# MonoDETRNext: Next-generation Accurate and Efficient Monocular 3D Object Detection Method

### v1

@@ 1 Introduction
3D object detection boasts a vast range of applications, from autonomous driving and robotic navigation to smart surveillance and virtual reality. These applications hinge critically on the precise recognition and localization of objects within 3D space. However, many 3D object detection methods often rely on expensive equipment such as LiDAR[1, 2, 3, 4, 5, 6], significantly limiting their practical deployment. While several purely visual strategies have achieved notable success, these generally involve multiple cameras[7, 8, 9] and are still limited under certain circumstances. Furthermore, they typically require substantial computational power to operate effectively.
Monocular 3D detection methods have not been given much attention, largely due to their relatively low accuracy. In many scenarios that require 3D detection, it is feasible to employ LiDAR or multiple cameras, supported by top-tier computing resources. However, in extreme cases where a LiDAR or one of the cameras fails, these high-precision algorithms might be rendered ineffective. Similarly, in scenarios with restricted device availability or budget constraints, multi-view and LiDAR-dependent algorithms may not be viable. Therefore, researching monocular 3D detection algorithms holds practical significance.
In the realm of 2D detection, models from the DETR series[10, 11, 12, 13] currently stand unmatched. MonoDETR[14] represents the first end-to-end monocular 3D object detection model within this category. Its performance surpasses all previous monocular 3D detection approaches. However, the impressive results of MonoDETR are not sufficient to satisfy our aspirations. Introducing some of the recent advancements in technology and strategies into monocular 3D vision detection, this paper presents a new-generation model, MonoDETRNext-F.
We also observed that in MonoDETR, the acquisition of depth information is overly simplistic, which is crucial for monocular 3D detection models since depth information is one of the few available 3D data types. The accuracy of depth information is closely linked to the precision of detection. To improve depth information acquisition, we referred to established 3D depth estimation networks and designed a depth predictor that directly extracts depth information from images. We integrated this predictor with MonoDETRNext-F to develop MonoDETRNext-A, which offers enhanced detection capabilities.
@@
    Proposing two novel monocular 3D object detection models, namely MonoDETRNext-F and MonoDETRNext-A, the former adeptly balancing speed and precision, while the latter accentuates precision-centric objectives.
@@
    Architecting a hybrid vision encoder for 3D object detection, proficient in the efficient extraction of features crucial for monocular 3D object detection tasks, alongside the integration of a refined, yet lightweight depth estimation module, envisaged to bolster detection accuracy.
@@
    Drawing upon methodologies hitherto entrenched in 2D object detection paradigms, we advance a cogent strategy for object query generation tailored to the exigencies of monocular 3D object detection, underpinned by meticulously designed training strategies, aimed at optimizing model performance efficaciously.
@@ 2 Related Work
Figure 1: Comparison of mainstream 3D detection models, with different colors representing distinct functional modules.
@@ 2.1 MonoDETR and other monocular 3D detection
Alternative methodologies, as demonstrated by MonoDLE [19], PGD [20], and PackNet [21], integrate multi-scale feature fusion and attention mechanisms for depth map estimation and error analysis, yielding improved performance. Despite their high accuracy, these approaches incur substantial computational costs and demand significant memory resources. Conversely, MonoDETR is characterized by its lightweight and efficient nature. Furthermore, MonoDETRNext-F surpasses it in speed and efficacy, while MonoDETRNext-A demonstrates markedly superior performance.
@@ 2.3 LiDAR and multi-source information fusion 3D object detection
Methods such as DeepFusion [30] and PointPainting [31] represent notable advancements in the integration of LiDAR point cloud data with camera imagery to facilitate precise object detection within three-dimensional spatial environments. This fusion strategy optimally exploits the synergies inherent in disparate sensor modalities, amalgamating spatial depth cues with color texture information, thereby fortifying the resilience and accuracy of detection outcomes.
The integration of principles from BEVFormer into fusion paradigms, exemplified by BevFusion [6], has spurred further refinements culminating in enhanced precision, as evidenced by exemplar models such as those delineated in MV2D [4] and Futr3d [32]. Recent endeavors, typified by mmFusion [33], have expanded the purview of fusion methodologies by integrating data from multiple sensors, including cameras, LiDAR, and radar, resulting in notable performance strides.
Concurrently, the domain has witnessed the emergence of large-scale architectures, exemplified by OMNI3D [34] and GLEE [35], which have showcased remarkable efficacy in 3D object detection tasks. Leveraging copious training data and sophisticated model architectures characterized by billions of parameters or more, these frameworks have been trained utilizing advanced optimization algorithms, thereby augmenting detection performance and precision.
@@ 3.1 Model Overview
The overall frameworks of MonoDETRNext-A and MonoDETRNext-F are illustrated in Figure 2. The main difference between these two models lies in the disparity of the depth predictors, which will be elaborated in Section 3.3. In Section 3.2, the design principles of the efficient encoder and the extraction of visual and depth features will be explicated. Lastly, in Section 3.4, the meticulous generation of object queries and the loss utilized will be expounded.
Figure 2: The schematic depiction of MonoDETRNext. The distinction between MonoDETRNext-A and MonoDETRNext-F primarily resides in their respective depth prediction mechanisms. The provided illustration delineates the intricate depth prediction scheme adopted by MonoDETRNext-A, whereas the depth predictor employed in MonoDETRNext-F remains congruent with that of MonoDETR.
3.2 Efficient Hybrid Vision Encoder
@@
Inspired by the architecture of RT-DETR[38], we introduce an innovative Efficient Hybrid Vision Encoder tailored specifically for 3d object detection tasks. This encoder is characterized by a diminished computational footprint while preserving the efficacy of feature extraction. Illustrated in Figure 2, our proposed encoder comprises two integral elements: a singular encoder layer and a Cross-Scale Feature Integration Module (CFIM) grounded in CNN. The CFIM functions as a fusion unit, amalgamating adjacent features into novel representations, as delineated in Figure 3. This fusion process is succinctly formalized by the ensuing equations:
[…]
@@
Compared to the hybrid encoder of RT-DETR, our hybrid encoder exhibits significant differences in the Fusion module.
3.3 Accurate Depth Predictor
@@
The principal disparity between our proposed models, MonoDETRNext-A and MonoDETRNext-F, resides in their Depth Predictor architecture. The initial Depth Predictor in MonodDETR, by design, featured a lightweight configuration comprising solely two $3$ convolutional layers, extracting $f_{D}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}\times(C)}$ from features preprocessed by the backbone. The accuracy of monocular 3D object detection correlates significantly with the quality of depth information obtained. In pursuit of refined depth estimation and subsequent enhancement of 3D detection precision, we undertook a redesign of the Depth Predictor. The revamped depth predictor now directly leverages the image data itself to extract depth-related features, circumventing secondary processing of backbone-derived features, thereby alleviating the potential for misalignment learning during model training.
@@
The architectural of this novel Depth Predictor draws inspiration from several state-of-the-art monocular depth estimation models, such as lite-mono[39] and monodepth2[40]. Its specific structural configuration, as depicted in Figure 2, initiates with the processing of images of dimensions $H\times W\times 3$ through a convolutional perturbation module, where image data undergoes downsampling via a $3\times 3$ convolutional operation. Subsequently, two additional $3\times 3$ convolutional layers, each with a stride of $1$, are deployed for local feature extraction, yielding feature maps of dimensions $\frac{H}{2}\times\frac{W}{2}\times C_{1}$. In the ensuing stage, these features are concatenated with the pooled three-channel input image, followed by another downsampling step through a $3\times 3$ convolutional layer with a stride of $2$, resulting in feature maps of dimensions $\frac{H}{4}\times\frac{W}{4}\times C_{2}$. The integration of features with the pooled input image in the downsampling layers serves to mitigate the loss of spatial information attributable to the reduction in feature dimensions, drawing inspiration from ESPNetv2[41]. Subsequent to this, RGFI and SDC are introduced to facilitate the acquisition of rich hierarchical feature representations. The ensuing downsampling layers also inherit connected features from the preceding downsampling layer. Analogously, the output feature maps undergo further processing until their dimensions align with those of MonoDETR[14], specifically $\frac{H}{16}\times\frac{W}{16}\times C_{4}$.
It is worth mentioning that MonoDETRNext-A utilizes our custom-designed Depth Predictor, while MonoDETRNext-F utilizes the original MonoDETR. After obtaining $f_{D}$, the subsequent operations for both our models, as well as MonoDETR, involve encoding through a linear classification layer and an encoder layer to obtain $D_{fg}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}\times(k+1)}$ along with other additional data. These data are then fed into the depth-guided decoder to parse features.
@@
The Sequential Dilated Convolution (SDC) module is introduced to extract multi-scale local features using dilated convolutions. Similar to lite-mono, we adopt a staged approach by inserting multiple consecutive dilated convolutions with varying dilation rates to effectively aggregate multi-scale context.
@@
$\hat{X}=X+\text{ Linear }_{G}\left(\text{ Linear }\left(BN\left(\text{ Conv }_{r}(X)\right)\right)\right),$ (6)
where $\text{Linear}_{G}$ denotes pointwise convolution followed by the GELU[42] activation. $BN$ represents the batch normalization layer, and $\text{Conv}_{r}(\cdot)$ is a $3\times 3$ convolutional layer.
@@
$\hat{X}=X+\text{ Linear }_{G}(\operatorname{Linear}(\operatorname{LN}(\tilde{X}))),$ (8)
where $LN$ represents a layer normalization[44] operation. Four variants of the depth encoder are designed based on different numbers of channels, SDC blocks. The complete process is illustrated in Figure 3 .
3.4 Effective Query Generation and Overall Loss
In DETR, object queries are a set of learnable embeddings optimized by the decoder and mapped to classification scores and bounding boxes via prediction heads. However, these object queries are challenging to interpret and optimize as they lack explicit physical meaning. Subsequent works[45, 46, 47, 11, 48] have improved the initialization of object queries and extended them to include content queries and position queries (anchors). In 3D object detection, the position of objects is no longer a simple 2D bounding box but includes depth information in a 3D bounding box. Therefore, a more accurate and interpretable method for generating 3D object queries is needed. To address this, we introduce the process of generating object queries from 2D DETR models into 3D object detection, expanding it into depth-aware object queries.
To put it simply, we have designed a target query generation method based on the characteristics of 3D objects. In MonoDETR[14], the final output of the bbox is a six-dimensional vector, $[x_{c},y_{c},x,y,w,h]$, $x_{c},y_{c}$ represent the 3D center point of the object in the image coordinates, and $x,y,w,h$ define the 2D bbox. However, MonoDETR’s object query generates only a two-dimensional vector, which is clearly not ideal. MV2D[4] provided us with inspiration, demonstrating that 2D bounding boxes can enhance the performance of 3D detection. Therefore, we incorporated the process of generating position queries from the 2D DETR model into our system, treating the queries as bounding boxes for loss computation. Specifically, the process is as follows:
where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}={\hat{c},\hat{b}}$ and $y={c,b}$, $c$ and $b$ represent categories and bounding boxes, respectively. We introduce the IoU score. Additionally, we incorporate the IoU score into the classification branch’s objective function (similar to VFL[49] and RT-DETR[38]), achieving consistency constraints for positive sample classification and localization.
For generating 3D centers, we experiment with several methods, but regretfully find that a set of learnable embeddings yields the best results. Finally, our overall loss is similar to that of MonoDETR, except $\mathcal{L}_{enc}$:
where $\mathcal{L}_{d\text{map }}$ represents the focal loss[50] of the predicted categorical foreground depth map $D_{fg}$ in Section.3.3 and $N_{gt}$ denotes the number of ground-truth objects. Additionally, it is noteworthy that the proportion of the total weight attributed to the $\mathcal{L}_{d\text{map}}$ loss during training differs between MonoDETRNext-A and MonoDETRNext-F, with MonoDETRNext-A having a higher percentage.

### v2

# MonoDETRNext: Next-Generation Accurate and Efficient Monocular 3D Object Detector

@@ Abstract
00footnotetext: Liao Pan, Yang Feng, Wu Di, Yu Jinwen, Zhao Wenhui are with Key Laboratory of Information Fusion, School of Automation, Northwestern Polytechnical University,Xi’an,China. (e-mile: liaopan@mail.nwpu.edu.cn,yangfeng@nwpu.edu.cn,wu_di821@mail. nwpu.edu.cn, yujinwen@mail.nwpu.edu.cn, zwh2024202513@mail.nwpu.edu.cn
@@ 1 Introduction
3D object detection boasts a vast range of applications, from autonomous driving and robotic navigation to smart surveillance and virtual reality. These applications hinge critically on the precise recognition and localization of objects within 3D space. However, many 3D object detection methods often rely on expensive equipment such as LiDAR [67, 15, 30, 66, 21, 34], significantly limiting their practical deployment. While several purely visual strategies have achieved notable success, these generally involve multiple cameras [33, 70, 51] However, these approaches are often multi-camera based, and they can be limited in certain situations, such as when a LiDAR or one or more cameras suddenly malfunction. Additionally, they generally require substantial computational power for support, which further reinforces their limitations. Monocular 3D detection methods, on the other hand, do not have these constraints. They can achieve satisfactory performance with just a single camera, making this class of methods highly valuable for research.
DETR-based models have demonstrated remarkable performance in both 2D detection [6, 84, 76, 9, 85] and multi-view [33, 70, 77, 64], multi-source 3D detection [12, 34, 5] . MonoDETR [79] was the first end-to-end monocular 3D object detection model in the DETR category. Upon its introduction, its performance surpassed all previous monocular 3D detection approaches. However, MonoDETR still exhibited certain limitations, with two notable issues: during query initialization, both the encoding and reference points lacked physical significance, and the depth estimation network was overly simplistic.
To address these issues, we first conducted an analysis of query generation for monocular 3D tasks and proposed a novel monocular 3D query generation method. This approach derives reference points and query encodings more effectively from depth and image features, significantly enhancing performance. Concurrently, we implemented a more efficient encoder to improve the model’s feature extraction capabilities while marginally increasing inference speed. These series of improvements led to the development of MonoDETRNext-E.
Subsequently, after experimenting with multiple established depth estimation networks, we identified a network that effectively enhanced MonoDETR’s detection performance. Based on this network and the aforementioned improvements, we introduced MonoDETRNext-A, a model with robust detection performance but slightly lower inference speed.
@@
    Proposing two novel monocular 3D object detection models, namely MonoDETRNext-E and MonoDETRNext-A, the former adeptly balancing speed and precision, while the latter accentuates precision-centric objectives.
@@
    Addressing the issue of nonsensical query initialization in MonoDETR, we introduced a bespoke object query generation strategy tailored for monocular 3D object detection, thereby refining model performance.
@@
    Evaluating numerous advanced networks, we identified two that were particularly effective. These findings facilitated the development of a new hybrid encoder and a streamlined, accurate depth estimation module for MonoDETR, significantly enhancing the model’s efficiency and detection precision.
@@ 2.1 MonoDETR and other monocular 3D detection
Alternative methodologies, as demonstrated by MonoDLE [48], PGD [63], and PackNet [49], integrate multi-scale feature fusion and attention mechanisms for depth map estimation and error analysis, yielding improved performance. Despite their high accuracy, these approaches incur substantial computational costs and demand significant memory resources. Conversely, MonoDETR is characterized by its lightweight and efficient nature. Furthermore, MonoDETRNext-E surpasses it in speed and efficacy, while MonoDETRNext-A demonstrates markedly superior performance.
@@ 2.2 Multi-view 3D object detection
Figure 1: Comparison of DETR-type 3D detection models, with different colors representing distinct functional modules.
@@ 2.3 LiDAR and multi-source information fusion 3D object detection
Methods such as DeepFusion [31] and PointPainting [60] represent notable advancements in the integration of LiDAR point cloud data with camera imagery to facilitate precise object detection within three-dimensional spatial environments. This fusion strategy optimally exploits the synergies inherent in disparate sensor modalities, amalgamating spatial depth cues with color texture information, thereby fortifying the resilience and accuracy of detection outcomes. The integration of principles from BEVFormer[33] into fusion paradigms, exemplified by BevFusion [34], has spurred further refinements culminating in enhanced precision, as evidenced by exemplar models such as those delineated in MV2D [66] and Futr3d [12]. Recent endeavors, typified by mmFusion [1], have expanded the purview of fusion methodologies by integrating data from multiple sensors, including cameras, LiDAR, and radar, resulting in notable performance strides.
Concurrently, the domain has witnessed the emergence of large-scale architectures, exemplified by OMNI3D [5] and GLEE [68], which have showcased remarkable efficacy in 3D object detection tasks.
@@ 3.1 Model Overview
Figure 2: The schematic depiction of MonoDETRNext. The distinction between MonoDETRNext-A and MonoDETRNext-E primarily resides in their respective depth prediction mechanisms. The provided illustration delineates the intricate depth prediction scheme adopted by MonoDETRNext-A, whereas the depth predictor employed in MonoDETRNext-E remains congruent with that of MonoDETR.
The overall frameworks of MonoDETRNext-A and MonoDETRNext-E are illustrated in Figure 2. A key focus of this paper is the meticulous generation of object queries, along with the associated loss, which will be expounded in Section 3.2.1. The main difference between these two models lies in the disparity of the depth predictors, which will be elaborated in Section 3.3.1. Additionally, the design principles of the efficient encoder and the extraction of visual and depth features will be explicated in Section 3.2.2.
3.2 MonoDETRNext-E
3.2.1 Depth-Guided Query Anchors
Figure 3: The differences in details between our decoder and MonoDETR’s decoder, the main difference between the two being the source of the embedding and the presence or absence of an initial anchors.
In Deformable-DETR [84], object queries are a set of learnable embeddings optimized by the decoder and mapped to classification scores and bounding boxes via prediction heads. However, these object queries are challenging to interpret and optimize as they lack explicit physical meaning. Subsequent works [28, 65, 37, 76, 73] have improved the initialization of object queries and extended them to include content queries and anchors. In 3D object detection, the position of objects is no longer a simple 2D bounding box but includes depth information in a 3D bounding box. Therefore, a more accurate and interpretable method for generating 3D object queries is needed. To address this, we introduce the process of generating object queries from 2D anchors into monocular 3D object detection, expanding it into depth-aware object queries.
To put it simply, we have designed a target query generation method based on the characteristics of 3D objects. In MonoDETR[79], the final output of the bbox is a six-dimensional vector, $[x_{c},y_{c},x,y,w,h]$, $x_{c},y_{c}$ represent the 3D center point of the object in the image coordinates, and $x,y,w,h$ define the 2D bbox. However, MonoDETR’s object query generates only a two-dimensional vector, consisting of a set of learnable embeddings, an approach that is clearly not ideal. MV2D[66] and QAF2D[24] provided us with inspiration, demonstrating that 2D bounding boxes can enhance the performance of 3D detection. Therefore, we incorporated the process of generating object queries from the 2D DETR model into our method, $x,y,w,h$ are directly learned from the image features, while the 3D center coordinates $x_{c},y_{c}$ are obtained by fusing depth information $f_{D}$ with visual features $f_{v}$. Specifically, the following formulations are used:
$&(x,y,w,h)=\text{MLP}_{1}(f_{v})\\ &(x_{c},y_{c})=\text{MLP}_{2}(\text{Attn}(f_{v},f_{D},f_{D}))\end{aligned}$ (1)
Here, $\text{MLP}_{1}$ is used to predict the 2D bounding box parameters solely from the visual features, while $\text{MLP}_{2}$ and Attn incorporate both the depth and visual features to estimate the 3D center coordinates. The inclusion of depth information in Attn allows the model to better capture 3D spatial relationships, leading to more accurate localization of the object’s 3D center.
One more point is that the query embedding we use is also different from MonoDETR, our embedding comes directly from the image features, the specific difference is shown in Figure 3.
3.2.2 Efficient Feature Extraction
It is well-established that efficient feature extraction methods often enhance model performance. In DETR-type models, feature extraction is primarily carried out by the backbone and encoder components. We experimented with various advanced networks for both components.
Regarding the backbone, essentially all our attempts were unsuccessful. Neither ConvNeXt [43], Swin Transformer [41], nor EfficientNet [54] outperformed ResNet [18] in effectiveness. However, our efforts with the Encoder proved fruitful. After modifying RT-DETR’s [82] Hybrid Encoder, we achieved performance improvements in our model. As illustrated in Figure 2, the enhanced encoder in this study comprises two main components: a single encoder layer and a Cross-Scale Feature Integration Module (CFIM). The integration module’s function is to combine adjacent features into a new feature representation, with its structure depicted in Figure 4. This fusion process is succinctly formalized by the ensuing equations:
@@ 3.2.2 Efficient Feature Extraction
Compared to the hybrid encoder of RT-DETR, our hybrid encoder exhibits differences in the Fusion module.
3.2.3 Overall Loss
In order to make the anchors work better, we also calculate the loss using the anchors obtained in the 3.2.1, and the $\mathcal{L}_{enc}$ is as follows
where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}={\hat{c},\hat{b}}$ and $y={c,b}$, $c$ represent categories and $b$ represent bounding boxes and 3D centers, respectively.
Finally, our overall loss is similar to that of MonoDETR, except $\mathcal{L}_{enc}$:
where $\mathcal{L}_{d\text{map }}$ represents the focal loss[35] of the predicted categorical foreground depth map $D_{fg}$ in Section.3.3.1 and $N_{gt}$ denotes the number of ground-truth objects. Additionally, it is noteworthy that the proportion of the total weight attributed to the $\mathcal{L}_{d\text{map}}$ loss during training differs between MonoDETRNext-A and MonoDETRNext-E, with MonoDETRNext-A having a higher percentage.
3.3 MonoDETRNext-A
@@
The principal disparity between our proposed models, MonoDETRNext-E and MonoDETRNext-A, resides in their Depth Predictor architecture. The initial Depth Predictor in MonodDETR, by design, featured a lightweight configuration comprising solely two $3$ convolutional layers, extracting $f_{D}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}\times(C)}$ from features preprocessed by the backbone. The accuracy of monocular 3D object detection correlates significantly with the quality of depth information obtained. In pursuit of refined depth estimation and subsequent enhancement of 3D detection precision, we undertook a redesign of the Depth Predictor. In order to further improve the accuracy of the model, we tried a variety of well-established depth estimation networks, and finally we chose the most effective one, the structure of this network we will show in this section.
3.3.1 Accurate Depth Predictor
Similar to the feature extraction process discussed in the previous section, we experimented with multiple state-of-the-art monocular depth estimation models, including Lite-Mono[78] , Depth-V2[17], and ZoeDepth [4]. Ultimately, only a simplified version of Lite-Mono improved detection performance. Conversely, other models led to a decline in detection accuracy. We posit that this outcome may be attributed to the challenge of balancing network training. It is crucial to note that our primary task is detection, not depth estimation. Excessively large depth estimation networks may potentially disrupt the training process, consequently leading to diminished detection performance.
@@ 3.3.1 Accurate Depth Predictor
The architectural of this Depth Predictor draws is depicted in Figure 2, initiates with the processing of images of dimensions $H\times W\times 3$ through a convolutional perturbation module, where image data undergoes downsampling via a $3\times 3$ convolutional operation. Subsequently, two additional $3\times 3$ convolutional layers, each with a stride of $1$, are deployed for local feature extraction, yielding feature maps of dimensions $\frac{H}{2}\times\frac{W}{2}\times C_{1}$. In the ensuing stage, these features are concatenated with the pooled three-channel input image, followed by another downsampling step through a $3\times 3$ convolutional layer with a stride of $2$, resulting in feature maps of dimensions $\frac{H}{4}\times\frac{W}{4}\times C_{2}$. , RGFI and SDC are introduced to facilitate the acquisition of rich hierarchical feature representations. The ensuing downsampling layers also inherit connected features from the preceding downsampling layer. Analogously, the output feature maps undergo further processing until their dimensions align with those of MonoDETR[79], specifically $\frac{H}{16}\times\frac{W}{16}\times C_{4}$.
@@
The Sequential Dilated Convolution (SDC) module is introduced to extract multi-scale local features using dilated convolutions. Similar to lite-mono, we adopt a staged approach by inserting multiple consecutive dilated convolutions with varying dilation rates to effectively aggregate multi-scale context. Compared to the CDC module of Lite-Mono, our SDC module has not point-wise layers, resulting in a more lightweight architecture.
@@
$\hat{X}=X+G\left(\left(BN\left(\text{ Conv }_{r}(X)\right)\right)\right),$ (9)
where $G$ denotes the GELU[20] activation. $BN$ represents the batch normalization layer, and $\text{Conv}_{r}(\cdot)$ is a $3\times 3$ convolutional layer.
@@
$\hat{X}=X+G(\operatorname{LN}(\tilde{X})),$ (11)
where $LN$ represents a layer normalization[3] operation. Similar to the SDC module, our RGF1 module does not include point-wise layers, unlike the LGF1 module of Lite-Mono [78]. The complete process is illustrated in Figure 4 .